A spacecraft takes 200 km after being launched to reach the escape speed from the earth of 11.2km/s. Find the average acceleration in terms of g

1
Expert's answer
2022-06-15T01:29:06-0400

Given:

"v=11.2\:\rm km/s=11200\: m/s"

"d=200\:\rm km=200000\: m"


"a=\frac{v^2}{2d}=\frac{11200^2}{2*200000}=313.6\:{\rm m/s^2}=32g"